OpenStudy (anonymous):

What is the solution to the following equation? x^2+14x+48=0 a)x=6 or x=8 b)x=-6 or x=-8 c)x=4 or x=12 d)x=-4 or x=-12

OpenStudy (anonymous):

@hero @hartnn @petiteme @tvartheman anyone?? :)

OpenStudy (jjuden):

b

OpenStudy (jjuden):

x^2+14x+48=0 Simple and best practice solution for x^2+14x+48=0 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for x^2+14x+48=0 equation: Simplifying x2 + 14x + 48 = 0 Reorder the terms: 48 + 14x + x2 = 0 Solving 48 + 14x + x2 = 0 Solving for variable 'x'. Factor a trinomial. (8 + x)(6 + x) = 0 Subproblem 1 Set the factor '(8 + x)' equal to zero and attempt to solve: Simplifying 8 + x = 0 Solving 8 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-8' to each side of the equation. 8 + -8 + x = 0 + -8 Combine like terms: 8 + -8 = 0 0 + x = 0 + -8 x = 0 + -8 Combine like terms: 0 + -8 = -8 x = -8 Simplifying x = -8 Subproblem 2 Set the factor '(6 + x)' equal to zero and attempt to solve: Simplifying 6 + x = 0 Solving 6 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-6' to each side of the equation. 6 + -6 + x = 0 + -6 Combine like terms: 6 + -6 = 0 0 + x = 0 + -6 x = 0 + -6 Combine like terms: 0 + -6 = -6 x = -6 Simplifying x = -6 Solution x = {-8, -6}

OpenStudy (anonymous):

oh wow thank you sooo much

OpenStudy (jjuden):

welcome :)